# Language Models as Hierarchy Encoders

### v1

@@ 4.2 Dataset Construction
Table 2: Multi-hop Inference and Mixed-hop Prediction test results on WordNet, comparing LMs of different sizes and different settings (pre-trained, fine-tuned, and hierarchy re-trained).
4.3 Baselines
Naive Prior
We first introduce a naive baseline that utilises the prior probability of positive subsumptions in the training set for prediction. Given that each positive sample is paired with $10$ negatives, the prior probability of a positive prediction stands at $\frac{1}{11}$. Consequently, Precision, Recall, and F-score on the test set are all $\frac{1}{11}$.
Pre-trained LMs
[…]
The pre-trained LMs selected from the sentence transformer library include all-MiniLM-L6-v2, all-MiniLM-L12-v2, and all-mpnet-base-v2, each re-trained on 1B sentence pairs from a spectrum of NLP tasks. [^12] https://discuss.huggingface.co/t/train-the-best-sentence-embedding-model-ever-with-1b-training-pairs/7354 The all-MiniLM models, derived from the MiniLM distilled from BERT (Wang et al., 2020), are noted for their efficiency and performance, with all-MiniLM-L6-v2 and all-MiniLM-L12-v2 being the variants of $6$ and $12$ layers, respectively. The embedding dimension of them is $384$. The all-mpnet-base-v2 model, based on MPNet (Song et al., 2020), is selected for its top performance in the sentence transformer collection131313 https://www.sbert.net/docs/pretrained_models.html. It has an embedding dimension of $768$. While the selected models are not specifically designed for hierarchy re-training, their optimisation for clustering and semantic search tasks makes them suitable initial checkpoints.
Fine-tuned LMs
Fine-tuned LMs are used as baselines to demonstrate that despite the efficacy in various tasks, standard fine-tuning struggles to address this specific challenge. Following the BERTSubs approach outlined in Chen et al. (2023), we employ pre-trained LMs with an added linear layer for binary classification, and optimising on the Softmax loss. Chen et al. (2023) have shown that this method outperforms various structure-based embedding techniques, such as TransE (Bordes et al., 2013) and DistMult (Yang et al., 2014), and also surpasses OWL2Vec* (Chen et al., 2021), which integrates both structural and textual embeddings, in the task of predicting subsumptions.
In Appendix B, we present additional comparison with static Poincaré embeddings (StaticPoincaré) (Nickel & Kiela, 2017) on the Multi-hop Inference task. We did not consider StaticPoincaré as a primary baseline because (i) this paper focuses on addressing limitations in LMs and (ii) StaticPoincaré cannot be applied to the Mixed-hop Prediction task (and the transfer settings) due to its fixed vocabulary for entities.
4.4 Implementation
The code implementation of this work primarily depends on DeepOnto (He et al., 2023a) for processing hierarchies and constructing datasets, Geoopt (Kochurov et al., 2020) for Poincaré ball, Sentence-Transformers (Reimers & Gurevych, 2019) and Huggingface-Transformers (Wolf et al., 2020) for training and evaluation of LMs. All our experiments were conducted on a single Quadro RTX 8000 GPU.
@@ 4.4 Implementation
In the hierarchy re-training of our HiT models, we configured the hyperbolic clustering loss margin ($\alpha$ in Equation 3) at $5.0$ and the hyperbolic centripetal loss margin ($\beta$ in Equation 4) at $0.1$. An exception was made for all-mpnet-base-v2 with hard negatives, where $\alpha$ was adjusted to $3.0$, based on validation. The models were trained for $20$ epochs, with a training batch size of $256$, $500$ warm-up steps and an initial learning rate of $10^{-5}$, using the AdamW optimiser (Loshchilov & Hutter, 2018). Model selection was conducted after each epoch, guided by performance on the validation set.
For fine-tuning, we largely adhered to the standard settings of the Huggingface Trainer [^14] https://huggingface.co/docs/transformers/main_classes/trainer, but maintained the same training batch size as used in the hierarchy re-training. Notably, our preliminary testing revealed that fine-tuning is more prone to overfitting. Consequently, we adopted a more frequent model selection interval, performing this assessment every $500$ training steps rather than epoch-wise.
4.5 Results
Performance of HiTs
The HiT models, re-trained from LMs of various sizes, consistently outperform their pre-trained and fine-tuned counterparts across all evaluation tasks. In the Multi-hop Inference task, HiTs exhibit exceptional performance with F-scores ranging from $0.871$ to $0.916$. This indicates a strong capability in generalising from asserted to transitively inferred entity subsumptions. In the Mixed-hop Prediction task, F-scores ranging from $0.856$ to $0.900$ highlight the effectiveness of HiTs in generalising from asserted to arbitrary entity subsumptions.
For the Transfer Mixed-hop Prediction tasks, we selected all-MiniLM-L12-v2 as the pre-trained model because all-MiniLM-L12-v2+HiT attains comparable performance to all-mpnet-base-v2+HiT while it is more computationally efficient owing to a smaller parameter size. Notably, all-MiniLM-L12-v2+HiT performs better than pre-trained and fine-tuned all-MiniLM-L12-v2 on these transfer tasks by at least $0.150$ and $0.101$ in F-scores, respectively.
@@ 4.5 Results
For the tasks on WordNet, all-mpnet-base-v2 achieves the highest F-scores among all the pre-trained models, yet these scores (e.g., $0.347$ and $0.250$ on the Mixed-hop Prediction task with random negatives and hard negatives, respectively) are considerably lower compared to their fine-tuned (lagging by $0.304$ and $0.419$) and hierarchy re-trained (lagging by $0.553$ and $0.613$) counterparts. This disparity confirms findings from LM probing studies such as those by Hanna & Mareček (2021) and He et al. (2023b), demonstrating the limited hierarchical knowledge in pre-trained LMs.
@@
The research by Chen et al. (2023) illustrates that fine-tuned LMs perform well on single-hop subsumptions. Our observations concur, showing that fine-tuned LMs achieve comparable performance as HiTs when assessed on just single-hop test samples. However, their effectiveness wanes when applied to arbitrary entity subsumptions. For the tasks on WordNet, fine-tuned LMs underperform HiTs by $0.194$ to $0.301$ in F-scores. In the transfer task from WordNet to Schema.org, the fine-tuned all-MiniLM-L12-v2 model only marginally outperforms its initial state, with an increase of around $0.02$ in F-scores across both negative settings.
@@
Figure 3: Visualisation of the distribution of WordNet entity embeddings generated by HiT (re-trained from all-MiniLM-L12-v2) w.r.t. their hyperbolic norms.
@@ 4.6 Analysis of HiT Embeddings
Figure 3 illustrates the distribution of WordNet entity embeddings generated by HiT w.r.t. their hyperbolic norms. These norms effectively capture the natural expansion of the hierarchical structure, evidenced by an exponential rise in the number of child entities. A notable observation is the sharp decline in the number of entities when hyperbolic norms exceed $23$, suggesting that few entities reside at these higher levels. Additionally, the range of entity hyperbolic norms, approximately from $8$ to $24$, indicates that a relatively small region of the high-dimensional manifold suffices to accommodate all entities in WordNet.
Case Study
@@ 5 Related Work
Prompt-based probing is widely used for extracting knowledge from LMs. Studies like Hanna & Mareček (2021) utilised cloze-style prompts for hypernym detection, while He et al. (2023b) approached subsumption prediction similar to Natural Language Inference. Lin & Ng (2022) examined if LMs, when correctly predicting “A is a B” and “B is a C”, can consistently infer the transitive relationship “A is a C”. These studies collectively highlight the limited capacity of pre-trained LMs in understanding hierarchical structures. Other research efforts, such as those by Liu et al. (2021) and Gosselin & Zouaq (2023), have aimed to incorporate structural semantics into LMs for entity encoding. However, these largely focus on entity equivalence or similarity, with less emphasis on hierarchical organisation.
Regarding hyperbolic embeddings, methods like the Poincaré embeddings (Nickel & Kiela, 2017) and the hyperbolic entailment cone (Ganea et al., 2018b) have shown effectiveness in representing hierarchies. However, these techniques are inherently static, limited by a fixed vocabulary size for entities, and lack inductive prediction capability – a limitation that can be mitigated by integrating them with LMs. Chen et al. (2020) ventured into this direction, applying learnable layers to project LM embeddings into hyperbolic space for syntax parsing and sentiment analysis. Our approach diverges from theirs by focusing on training LMs as general hierarchy encoders without the need for additional learnable parameters.
@@ Appendix A Hierarchy Construction from Ontologies
The Terminology Box (TBox) in an OWL (Web Ontology Language) [^17] https://www.w3.org/OWL/ ontology defines entity relationships using subsumption axioms ($C\sqsubseteq D$) and equivalence axioms ($C\equiv D$), where $C$ and $D$ are atomic or complex concept expressions defined in the description logic $\mathcal{SROIQ}$. In this work, we used atomic concepts as nodes of the hierarchy. We employed an ontology reasoner to deduce direct [^18] Refer to the definition of $DirectSubClassOf$ at https://owlcs.github.io/owlapi/apidocs_4/org/semanticweb/owlapi/reasoner/OWLReasoner.html. subsumptions and included these as edges in the hierarchy. This approach fully considers both subsumption and equivalence axioms for all potential edges. Considering an atomic concept $Beef$, which is defined by $Beef\equiv Meat\sqcap\exists derivesFrom.Cattle$, as an example, the reasoning will lead to an edge between $Beef$ and $Meat$. Without reasoning, $Beef$ will be misplaced under the root node, if no other subsumptions about $Beef$ are asserted in the ontology. Tools like Protégé [^19] https://protege.stanford.edu/ also demonstrate similar considerations in presenting ontology concept taxonomies.
@@
It is also important to note the variability in naming schemes across different ontologies, which sometimes necessitates pre-processing of entity names. He et al. (2023b) provided detailed pre-processing steps for Schema.org, FoodOn, and DOID. In this study, we used the pre-processed versions of FoodOn and DOID and applied the same pre-processing methodology to the latest version of Schema.org.
Appendix B Results for Static Poincaré Embeddings
Random Negatives Hard Negatives
Model Precision Recall F-score Precision Recall F-score
Multi-hop Inference (WordNet)
StaticPoincaré 0.862 0.866 0.864 0.797 0.867 0.830
all-MiniLM-L12-v2 0.127 0.585 0.209 0.108 0.740 0.188
+ fine-tune 0.811 0.515 0.630 0.819 0.530 0.643
+ HiT 0.880 0.927 0.903 0.910 0.906 0.908
Table 5: Multi-hop Inference test results on WordNet, comparing StaticPoincaré and the HiT model re-trained from all-MiniLM-L12-v2.
Table 5 provides additional comparison to the static Poincaré embeddings (StaticPoincaré) in the Multi-hop Inference task on WordNet. Although StaticPoincaré surpasses the pre-trained and fine-tuned all-MiniLM-L12-v2 in terms of F-scores, it still falls short of all-MiniLM-L12-v2+HiT, by margins of $0.039$ and $0.105$ in the random and hard negative settings, respectively. This comparison further highlights the effectiveness of integrating hyperbolic geometry with language models.
A crucial advantage of HiT over StaticPoincaré is its capability for inductive prediction, enabling its application in tasks like Mixed-hop Prediction and transfer learning scenarios, where StaticPoincaré is not applicable due to its static vocabulary for entities.
Notably, the random negative setting in the Multi-hop Inference task follows the evaluation framework in Ganea et al. (2018b). Our implementation of StaticPoincaré not only outperforms their baseline models but also shows better results than their hyperbolic entailment cone model, even though the latter was trained with an extra $10\%$ of multi-hop edges.
Appendix C Results on SNOMED CT
@@
In the main body of this paper, we primarily focused on models trained on the WordNet (Noun). This section broadens our study to encompass models trained on SNOMED CT (Stearns et al., 2001), a structured, comprehensive, and widely-used vocabulary for electronic health records. Using the tool provided by the SNOMED CT team, [^20] https://github.com/IHTSDO/snomed-owl-toolkit we converted the latest version of SNOMED CT (released in December 2023) into the OWL ontology format. We then constructed the hierarchy according to the procedure detailed in Appendix A.
@@
Table 7: Mixed-hop Prediction test results on SNOMED, comparing LMs of different sizes and different settings (pre-trained, fine-tuned, and hierarchy re-trained), and Transfer Mixed-hop Prediction results on Schema.org, FoodOn, and DOID.

### v2

@@ 4.3 Baselines
4.4 Results
Performance of HiTs
The HiT models, re-trained from LMs of various sizes, consistently outperform their pre-trained and standard fine-tuned counterparts across all evaluation tasks. In the Multi-hop Inference task, HiTs exhibit exceptional performance with F-scores ranging from $0.871$ to $0.916$. This indicates a strong capability in generalising from asserted to transitively inferred entity subsumptions. In the Mixed-hop Prediction task, F-scores ranging from $0.856$ to $0.900$ highlight the effectiveness of HiTs in generalising from asserted to arbitrary entity subsumptions. For the Transfer Mixed-hop Prediction tasks, we selected all-MiniLM-L12-v2 as the pre-trained model because all-MiniLM-L12-v2+HiT attains comparable performance to all-mpnet-base-v2+HiT while it is more computationally efficient owing to a smaller parameter size. Notably, all-MiniLM-L12-v2+HiT performs better than pre-trained and fine-tuned all-MiniLM-L12-v2 on these transfer tasks by at least $0.150$ and $0.101$ in F-scores, respectively.
@@ 4.4 Results
For the tasks on WordNet, all-mpnet-base-v2 achieves the highest F-scores among all the pre-trained models, yet these scores (e.g., $0.347$ and $0.250$ on the Mixed-hop Prediction task with random negatives and hard negatives, respectively) are considerably lower compared to their fine-tuned (lagging by $0.304$ and $0.419$) and hierarchy re-trained (lagging by $0.553$ and $0.613$) counterparts. This disparity confirms findings from LM probing studies such as those by [5] and [6], demonstrating the limited hierarchical knowledge in pre-trained LMs.
@@
The research by [8] illustrates that fine-tuned LMs perform well on single-hop subsumptions. Our observations concur, showing that fine-tuned LMs achieve comparable performance as HiTs when assessed on just single-hop test samples. However, their effectiveness wanes when applied to arbitrary entity subsumptions. For the tasks on WordNet, fine-tuned LMs underperform HiTs by $0.194$ to $0.301$ in F-scores. In the transfer task from WordNet to Schema.org, the fine-tuned all-MiniLM-L12-v2 model only marginally outperforms its initial state, with an increase of around $0.02$ in F-scores across both negative settings.
Performance of Hyperbolic Baselines
The Multi-hop Inference task with random negatives follows the evaluation in [13]. In this setup, both PoincaréEmbed and HyperbolicCone significantly outperform the pre-trained and standard fine-tuned LMs, and perform comparably to the HiT models. However, HyperbolicCone exhibits substantially worse performance in the hard negative setting; its low precision and high recall suggest difficulties in differentiating sibling entities that are closely positioned in the embedding space. In the transfer evaluation, PoincaréGloVe shows improved performance over pre-trained and standard fine-tuned models on Schema.org. However, it does not demonstrate a similar advantage on FoodOn and DOID, primarily due to its limited vocabulary, which allows it to predict almost all test samples on Schema.org but substantially fewer on the others.
@@
More discussion on loss functions and an ablation study of loss margins are presented in Appendix C.
@@ 4.5 Analysis of HiT Embeddings
Figure 3 illustrates how WordNet entity embeddings generated by all-MiniLM-L12-v2+HiT distribute w.r.t. their hyperbolic norms.
Figure 3: Distribution of WordNet entity embeddings generated by HiT w.r.t. their hyperbolic norms.
Table 4: Statistical correlations between WordNet entities’ depths and their hyperbolic norms across different hyperbolic models.
HiT PoincaréEmbed HyperbolicCone
0.346 0.130 0.245
@@
These norms effectively capture the natural expansion of the hierarchical structure, evidenced by an exponential rise in the number of child entities. A notable observation is the sharp decline in the number of entities when hyperbolic norms exceed $23$, suggesting that few entities reside at these higher levels. Additionally, the range of entity hyperbolic norms, approximately from $8$ to $24$, indicates that a relatively small region of the high-dimensional manifold suffices to accommodate all entities in WordNet.
Correlation
In Table 4, we compare the Pearson correlation coefficients across different hyperbolic models to measure the linear relationship between entities’ hyperbolic norms and their depths in WordNet. Our analysis shows that all hyperbolic models lead to a positive correlation between norms and depths, as expected. However, HiT demonstrates a stronger correlation than both PoincaréEmbed and HyperbolicCone.
Case Study
@@ 5 Related Work
Prompt-based probing is widely used for extracting knowledge from LMs. Studies like [5] utilised cloze-style prompts for hypernym detection, while [6] approached subsumption prediction similar to Natural Language Inference. [7] examined if LMs, when correctly predicting “A is a B” and “B is a C”, can consistently infer the transitive relationship “A is a C”. These studies collectively highlight the limited capacity of pre-trained LMs in understanding hierarchical structures. Other research efforts, such as those by [10] and [11], have aimed to incorporate structural semantics into LMs for entity encoding. However, these largely focus on entity equivalence or similarity, with less emphasis on hierarchical organisation.
Regarding hyperbolic embeddings, methods like the Poincaré embedding [12] and the hyperbolic entailment cone [13] have effectively represented hierarchical structures. Despite their efficacy, these techniques are inherently static, constrained by a fixed vocabulary of entities, and do not support inductive predictions about unseen data. Further explorations include learning word embeddings in hyperbolic space [31, 32]. These methods, however, are limited to word-level tokenisation and yield non-contextual word representations. These shortcomings can be mitigated by integrating hyperbolic embeddings with transformer-based LMs. [33] has explored this direction, applying learnable layers to project LM embeddings into hyperbolic space for syntax parsing and sentiment analysis. Our approach diverges from theirs by focusing on training LMs as general hierarchy encoders without the need for additional learnable parameters.
@@ 6 Conclusion
Limitations
This work does not address the potential loss of pre-trained language understanding resulted from hierarchy re-training. Also, the issue of entity naming ambiguity inherent in the dataset sources is not tackled, which could introduce noise into the training process.
Acknowledgments and Disclosure of Funding
This work was supported by Samsung Research UK (SRUK), and EPSRC projects OASIS (EP/S032347/1), UK FIRES (EP/S019111/1), and ConCur (EP/V050869/1). Special thanks to Zifeng Ding for his valuable feedback during the rebuttal process.
@@ Appendix A Hierarchy Construction from Ontologies
The Terminology Box (TBox) in an OWL (Web Ontology Language) [^12] https://www.w3.org/OWL/ ontology defines entity relationships using subsumption axioms ($C\sqsubseteq D$) and equivalence axioms ($C\equiv D$), where $C$ and $D$ are atomic or complex concept expressions defined in the description logic $\mathcal{SROIQ}$. In this work, we used atomic concepts as nodes of the hierarchy. We employed an ontology reasoner to deduce direct [^13] Refer to the definition of $\mathsf{DirectSubClassOf}$ at https://owlcs.github.io/owlapi/apidocs_4/org/semanticweb/owlapi/reasoner/OWLReasoner.html. subsumptions and included these as edges in the hierarchy. This approach fully considers both subsumption and equivalence axioms for all potential edges. Considering an atomic concept $\mathsf{Beef}$, which is defined by $\mathsf{Beef\equiv Meat\sqcap\exists derivesFrom.Cattle}$, as an example, the reasoning will lead to an edge between $\mathsf{Beef}$ and $\mathsf{Meat}$. Without reasoning, $\mathsf{Beef}$ will be misplaced under the root node, if no other subsumptions about $\mathsf{Beef}$ are asserted in the ontology. Tools like Protégé [^14] https://protege.stanford.edu/ also demonstrate similar considerations in presenting ontology concept taxonomies.
@@
It is also important to note the variability in naming schemes across different ontologies, which sometimes necessitates pre-processing of entity names. [6] provided detailed pre-processing steps for Schema.org, FoodOn, and DOID. In this study, we used the pre-processed versions of FoodOn and DOID and applied the same pre-processing methodology to the latest version of Schema.org.
Appendix B Experiment Settings
The code implementation of this work primarily depends on DeepOnto [34] for processing hierarchies and constructing datasets, Geoopt [35] for Poincaré ball, Sentence-Transformers [21] and Huggingface-Transformers [36] for training and evaluation of LMs. All our experiments were conducted on a single Quadro RTX 8000 GPU.
In the hierarchy re-training of our HiT models, we configured the hyperbolic clustering loss margin ($\alpha$ in Equation 3) at $5.0$ and the hyperbolic centripetal loss margin ($\beta$ in Equation 4) at $0.1$. An exception was made for all-mpnet-base-v2 with hard negatives, where $\alpha$ was adjusted to $3.0$, based on validation. The models were trained for $20$ epochs, with a training batch size of $256$, $500$ warm-up steps and an initial learning rate of $10^{-5}$, using the AdamW optimiser [37]. Model selection was conducted after each epoch, guided by performance on the validation set.
For standard fine-tuning, we largely adhered to the default settings of the Huggingface Trainer [^15] https://huggingface.co/docs/transformers/main_classes/trainer, but maintained the same training batch size as used in the hierarchy re-training. Notably, our preliminary testing revealed that fine-tuning is more prone to overfitting. Consequently, we adopted a more frequent model selection interval, performing this assessment every $500$ training steps rather than epoch-wise.
For our hyperbolic baselines, we trained PoincaréEmbed with an embedding dimension of $200$ for $200$ epochs, a training batch size of $256$, $10$ warm-up epochs, and a constant learning rate of $0.01$, using the Riemannian Adam optimiser [38]. According to [13], HyperbolicCone benefits from a robust initialisation such as the one provided by a pre-trained PoincaréEmbed. Consequently, we utilised the same hyperparameters to train HyperbolicCone except that we initialised it with the weights from our pre-trained PoincaréEmbed. As outlined in the main paper, we selected the optimal pre-trained PoincaréGloVe model reported by [32] (50$\times$2D with an initial trick) as a baseline for our transfer evaluation.
Appendix C Further Discussion on $\mathcal{L}_{\textsc{HiT}}$
Loss Variants
As discussed in the main paper, we opted for the triplet contrastive loss format for hierarchy re-training because hierarchically related entities (e.g., subsumptions) should be closer together, yet not equivalent. We tested other forms of loss, including standard contrastive loss, which minimises absolute hyperbolic distances between related entities, and softmax contrastive loss, which was adopted in training PoincaréEmbed [12]. Our trial experiments demonstrated that the triplet form converges more efficiently and effectively compared to these alternatives.
Loss Margins
We provide an ablation study on loss margins $\alpha$ and $\beta$ defined in Equation (3) and Equation (4), respectively.
Table 6: Ablation results (F-score) of allMiniLM-L12-v2+HiT on WordNet’s Mixed-hop Prediction.
$\alpha=5.0,\beta=0.1$ $\alpha=3.0,\beta=0.1$ $\alpha=1.0,\beta=0.1$ $\alpha=5.0,\beta=0.5$
0.885 0.865 0.867 0.899
The results from Table 6 indicate that although loss margins impact performance, the HiT model exhibits robustness to their variations. Notably, a higher F-score for $\alpha=5.0,\beta=0.5$ is observed, surpassing the results presented in the main paper. Despite this, we chose not to overly optimise $\alpha$ and $\beta$ to avoid overfitting and to maintain the generalisability of our findings.
Appendix D Results on SNOMED CT
@@
In the main paper, we primarily focused on models trained on the WordNet (Noun). This section broadens our study to encompass models trained on SNOMED CT [15], a structured, comprehensive, and widely-used vocabulary for electronic health records. Using the tool provided by the SNOMED CT team, [^16] https://github.com/IHTSDO/snomed-owl-toolkit we converted the latest version of SNOMED CT (released in December 2023) into the OWL ontology format. We then constructed the hierarchy according to the procedure detailed in Appendix A.
@@
Table 8: Mixed-hop Prediction test results on SNOMED and Transfer Mixed-hop Prediction results on Schema.org, FoodOn, and DOID.